A balloon is ascending at the rate of 10 m/s2 at a height 40 m above the ground when a packet is dropped.How long will it take for the packet to reach the ground?

Grade:11

1 Answers

Ayush Jain
6 years ago
As the ballon was intially going up soU =-10m/s and a=g,s=40m/s SoS=ut+1/2at^240=-10t+5t^2Divided by 58=-2t+t^2T^2-2+-8=0(t-4)(t+2)=0Hence as time can`t be negative so time =4s